Show that if A and B are sets with the same cardinality, then |A|<=|B| and |B|<=|A|
1
Expert's answer
2020-10-22T18:04:58-0400

We say that "|X|\leq |Y|" if there is an injection "i: X\to Y". If the sets "A" and "B" are of the same cardinality, then there exists a bijection "f:A\to B". Since each bijection is an injection, "f" is an injection, and therefore "|A|\leq |B|". The inverse function "f^{-1}:B\to A" is a bijection as well. Consequently, "f^{-1}" is an injection. We conclude that "|B|\leq |A|", and we are done.